# IncomeSCM: From tabular data set to time-series simulator and causal estimation benchmark

### v2

@@ 1 Introduction
Simulators can be data-driven, learned from observational as in Medkit-Learn (Chan et al., 2021), or hand-designed as in the 2016 ACIC causal inference challenge (Dorie et al., 2019). Hand-designed simulators are great for designing benchmarks to test estimators in a particular challenge, such as learning from non-linear outcomes (Hill, 2011; Dorie et al., 2019) or heteroskedastic noise (Hahn et al., 2019). Their main drawbacks are that they i) scale poorly with the size of the system, requiring significant domain expertise by the designer, or that they ii) model overly simple or extreme systems, not reflecting the intricacies of real-world data (Kinyanjui and Johansson, 2022; Hernán, 2019). In contrast, purely data-driven simulators of potential outcomes can fit and generate complex relations between context, treatment, and outcome variables which are difficult to design by hand (Chan et al., 2021). However, they often lack structural constraints on the causal relationships between variables. As a result, they are insufficient for reasoning about different identification strategies, e.g., with instrumental variables, or with different adjustment sets, and for changing the nature of confounding or interventions (Hernán et al., 2019). In other words, mimicing only the observational distribution of a chosen system is not sufficient to stress test causal effect estimators.
@@
In this work, we construct a simulator using a sequential structural causal model based on the well-known Adult data set (Becker and Kohavi, 1996) and use this to create a new benchmark data set for causal effect estimation. We apply a generalizable strategy for developing realistic and challenging estimation tasks by imposing structural constraints on the simulator using a causal graph and fitting parts of its mechanisms to non-sequential tabular data. Our strategy assumes that the initial state of a system of variables follows a complex distribution (e.g., the age, education and income of a subject), but that its transitions are simpler (e.g., the income next year is close to the income from the previous year). Thus, we can build a realistic simulator by fitting the causal mechanisms of initial states to data, and hand-designing transition mechanisms. The effects of interventions in early time steps on outcomes later in time are determined by compositions of several transitions which renders causal effects non-trivial functions of the initial state, even if each transition is simple. We use our simulator IncomeSCM to generate data for an observational study of the effects of studies on future income and compare popular estimators from the literature. Our results indicate that different estimators yield estimates that differ substantially in quality despite fitting observed variables similarly well, pointing to the challenging nature of the tasks.
@@ 2 Related work
Recognizing the limitations of relying on purely synthetic simulators for benchmarking, several works have proposed fully or partially data-driven simulators, where components are learned from real-world data. A recent example is Medkit-Learn Chan et al. (2021), developed for off-policy evaluation of reinforcement learning, and the Alzheimer’s Disease Causal estimation Benchmark (ADCB) (Kinyanjui and Johansson, 2022). The 2022 iteration of the ACIC challenge included a partially data-driven design, which used a strategy similar to ours, but to the best of our knowledge, the details of the simulator remain unpublished (Mathematica, 2022). Complementing purpose-built data sets and simulators, causality toolboxes, such as DoWhy (Sharma and Kiciman, 2020) and CausalML (Chen et al., 2020) include multiple ways of generating simple forms of synthetic data. Gentzel et al. (2019) provide an excellent overview of different approaches to evaluating causal modeling methods and suggest another approach: Use complex systems, such as a PostGres data base, where interventions on its configuration can be performed but the outcome are hard to predict. This can provide both observational data for learning and interventional data for evaluation.
@@ 3 Constructing the simulator
IncomeSCM is based on a discrete-time Markov structural causal model (SCM). The primary goal of the model is to simulate the effects of an intervention of interest $A$ on a given outcome $Y$ in a system of random variables $\mathcal{V}\coloneqq\{A,Y\}\cup{\bf X}$, where ${\bf X}$ is a set of context variables. Each variable $V\in\mathcal{V}$ is simulated at $T$ time points in a sequence $V_{1},...,V_{T}$. The initial state of $V_{1}$ is determined by a deterministic structural equation $V_{1}:=f_{V}(\mathrm{\textbf{pa}}(V_{1}),U_{V_{1}})$ where $\mathrm{\textbf{pa}}(V_{1})$ are the parents (direct causes) of $V_{1}$ and $U_{V_{1}}$ is a source of noise. Values of $V_{t}$ for $t>1$ are determined by a time-homogenous transition mechanism $V_{t}:=g_{V}(\mathrm{\textbf{pa}}(V_{t}),U_{V_{t}})$. In general, mechanisms and parent sets are not shared by initial states and transitions, i.e., $f_{V}\neq g_{V}$ and $\mathrm{\textbf{pa}}(V_{1})\neq\mathrm{\textbf{pa}}(V_{t})=\mathrm{\textbf{pa}}(V_{s})$ for $s,t>1$. The model is Markov in that $\mathrm{\textbf{pa}}(V_{t})$ includes only variables from time $t$ and $t-1$.
@@ 3.1 The base data set: Adult
As primary outcome of interest $Y$, we use the yearly income of subjects in US dollars. In the Adult data set, income is coded as a binary variable, representing whether the yearly income of a subject exceeds \$50 000, denoted here by $R\in\{0,1\}$. For IncomeSCM, we construct a continuous income variable $Y$ by first fitting a small random forest regressor $h_{R}$ to predict the original binary outcome $R$ as a function of all covariates $X$ (excluding the added variable studies). The continuous predictions $\hat{R}_{i}=h_{R}(X_{i})\in[0,1]$ of the regressor are computed for each subject $i$ and re-scaled by a constant $c$ to achieve a desired average income across the full cohort.
@@
For a causal estimation task to be realistic, the primary intervention of interest must be possible to implement in practice, for example, in an experiment (Hernán, 2019; Hernán et al., 2022). The Adult data set contains a small number of variables that could be directly intervened upon and serve as the primary intervention. Base variables age, sex, native-country, and race can be removed from consideration, as can capital-gain, capital-loss as these are outside of the control of any feasible experiment. occupation is feasible to change on a yearly basis, but most jobs are unavailable to most people. relationship and marital-status are possible to change, through e.g., divorce, but most people would not be willing to change these in order to optimize their income, let alone participate in an experiment to do so. workclass, occupation and hours-per-week are candidates that could be intervened upon a yearly basis. education is coded in Adult by the highest attained degree (e.g., Bachelor’s degree) or the number of years of study. However, in the time-scale of the simulator (one observation per year, $t=1,...,T$), an intervention of, say $A_{t}\leftarrow\mathrm{Doctorate}$ is not feasible. Instead, to showcase the blend of data-driven and knowledge-guided design, we create an intervention on studies, a constructed variable, with values “Full-time studies”, “Day course”, “Evening course” and “No studies”, representing the study activity of a person during a year. The initial state of studies is determined by a logistic regression with manually selected coefficients for age, relationship, and education, and intercepts for each of the study types. studies directly affect the education level, current income and the propensity to study in the next year, and indirectly the future income of a subject.
@@ 3.2 The causal graph
We construct a cDAG for the Adult data set (the initial-state cDAG for IncomeSCM-1.0) by assigning direct causes to each variable. age, sex, race and native-country were taken to be base variables without causal parents in the data set. education was deemed to be causal of workclass, which in turn causes occupation and marital-status. marital-status and relationship are closely linked, and the causal direction was chosen to be from the former to the latter. Finally, hours-per-week and capital-net were determined by the previous variables, and income is the furthest downstream, directly caused also by studies, the hand-designed variable. The full graph is illustrated in Figure 1 (middle-left). A very similar graph was used by Nabi and Shpitser (2018) except, in their case, marital-status was a direct cause of education, not the other way around.
@@ 3.3 The samplers
where $f_{V}:\mathrm{\textbf{pa}}(V)\rightarrow\mathbb{R}$ predicts $V$ given its parents—$f_{V}$ is a model of the conditional mean, $\mathbb{E}[V\mid\mathrm{\textbf{pa}}(V)]$—and $\epsilon_{V}$ is a source of homoskedastic noise (e.g., Gaussian). All structural equations for initial-state variables, except for studies, have observations $(V,\mathrm{\textbf{pa}}(V))$ available (income is a special case as described above). For these, $f_{V}$ is implemented by fitting a regression (e.g., a Random forest regressor) to the observed data, and the scale of $\epsilon_{V}$ is chosen proportional to the mean squared regression error. capital-net is an exception to this form, since the majority of capital gains/losses is 0. For this, we blend a classifier which predicts non-zero values and a conditional regressor for the case where the value is non-zero.
For categorical variables $V\in\{1,...,k\}$, we sample from a point on the $k$-simplex $\Delta_{k}$ given by a map $f_{V}:\mathrm{\textbf{pa}}(V)\rightarrow\Delta_{k}$. By the Gumbel-Max trick (Gumbel, 1954), this is equivalent to
@@
For equations with no available data, such as for studies, $f_{V}$ is designed by hand. For studies, both the initial-state distribution and transitions are hand-crafted, as described above. For income1 and the transitions $g_{V}$ of most variables $V\in\mathcal{V}$, we use a blend of hand-designed rules and fitted regressions. This allows us to generate complex dependencies between variables without observing a time-series. For example, income1 is based on a regression of the original binary variable in the Adult data set, with hand-crafted modifications depending on studies and workclass. The transition sampler of relationshipt has a fixed probability $p_{\mathrm{stay}}$, that the person will stay in their relationship, relationshipt = relationshipt-1. With probability $1-p_{\mathrm{stay}}$, relationshipt is resampled from the same fitted model as the initial state, see Figure 1 (right). In this way, only the event that the relationship is changed is hand-designed, but what it is changed to depends on data.
@@ 4 Implementation & data set
Here, parents specify the direct causes of the initial state of the variable. sampler specifies the type and parameters of the sampler. In this case, the initial state of education is distributed according to a logistic regression of age, race, sex and native-country. seq_parents_curr specifies the direct causes of the transition probability from the current time step, and seq_parents_curr causes from the previous time step. seq_sampler specifies which function determines this probability. LogisticSampler is a general sampler, that outputs the probability of the given variable as a logistic-linear function of the direct causes specified in parents. General samplers can be reused for other data sets and are defined in samplers.py.
@@ 4.1 Benchmark data set
For this task, we create a cross-sectional observational data set by first sampling from the simulator at time $t=1,...,7$ where studies are selected by a “default” observational stochastic policy as defined by the samplers in IncomeSCM. Then, time $t=1$ is used to record the previous income level income_prev and studies studies_prev, prior to intervention. All other covariates are read off at time $t_{0}=2$. The binary intervention of interest is then performed at $t_{0}=2$, with $A_{2}=1$ representing Full-time studies and $A_{2}=0$ No studies. The outcome variable $Y_{T}=$incomeT is recorded at $T=7$. We use this to sample $n=50000$ samples with random seed $s=0$. First-order covariate statistics of covariates, interventions and outcomes for this seed can be seen in Table 1.
@@
    CATE, conditioned only on education (not itself a valid adjustment set), represented as 16 numeric bins, using the full adjustment set from Task 1.
Task 3 is an example of when the conditioning set differs from any valid adjustment set, requiring a slightly unusual application of common causal effect estimators.
Valid adjustment sets can be derived by standard graphical criteria like the backdoor criterion (Pearl, 2009) from the edge list in Table 5. Note, as described above, that the observed intervention is performed at time $t=2$ with $t=1$ being the initial state. Thus, a minimal adjustment set for estimating the causal effect of intervening on studies2 is the set of direct causes of this variable. In IncomeSCM-1.0, studiest at times $t>0$ are directly caused by studiest-1, incomet-1, aget, sext,educationt, and relationshipt. [^2] In the implementation, the helper variable time is also passed to enable interventions at specific time steps. Expanding this set with any pre-intervention variables (such as other variables at times $t=1,2$) also results in valid adjustment sets. However, to estimate the CATE conditioned only on education2, it is not sufficient to adjust only for this variable.
@@ 5 Experiments
We include regression estimators in the form of T-learners, which fit one model per potential outcome, and S-learners, which treat the intervention as an input to a single model, mixed in with adjustment variables (Künzel et al., 2019). For both, we compare multiple base estimators in ridge regression (Ridge), random forests (RF), and XGBoost (XGB). Second, we use two classical inverse-propensity weighting estimators, the Horvitz-Thompson (IPW) and Hayek or “Weighted” (IPW-W). Finally, we include a nearest-neighbor matching estimator with a Euclidean metric. The IPW, matching and Ridge-regression S-learner estimators return estimates only of the ATE, not the CATE. For a more detailed description of all estimators, see Appendix E.
Hyperparameters for all estimators (see Appendix Table 6) were selected based on 5-fold cross-validation of 20 uniformly sampled hyperparameter settings. The estimators were then refitted to the entirety of the observational training data. Producing all results, including fitting the simulator, sampling, estimating and evaluating causal effects, took less than 1 hour on a 2023 M2 MacBook Pro. The results were replicated on a 2 x 16 core Intel(R) Xeon(R) Gold 6226R CPU @ 2.90GHz with minimal differences except for XGBoost which is known to yield different fits for different machines.

### v3

@@ 1 Introduction
Hand-designed simulators are great for designing benchmarks to test estimators in a particular challenge, such as learning from non-linear outcomes (Hill, 2011) or heteroskedastic noise (Hahn et al., 2019). Their main drawbacks are that they (i) scale poorly with the size of the system, requiring significant domain expertise by the designer, or (ii) model overly simple or extreme systems, not reflecting the intricacies of real-world data (Hernán, 2019). In contrast, purely data-driven simulators of potential outcomes can fit and generate complex relations between context, treatment, and outcome variables which are difficult to design by hand (Chan et al., 2021). However, they often lack structural constraints on the causality between variables, making them insufficient for reasoning about rivaling identification strategies, e.g., with instrumental variables or alternative adjustment sets, and for changing the nature of confounding or interventions (Hernán et al., 2019). In short, mimicking only the observational distribution of a chosen system is not sufficient to stress test causal effect estimators.
@@
In this work, we construct a simulator using a sequential structural causal model based on the well-known Adult data set (Becker and Kohavi, 1996) and use this to create a new benchmark data set for causal effect estimation. We apply a generalizable strategy for developing realistic and challenging estimation tasks by imposing structural constraints on the simulator using a causal graph and fitting parts of its mechanisms to non-sequential tabular data. Our strategy assumes that the initial state of a system of variables follows a complex distribution (e.g., the age, education, and income of a subject), but that its transitions are simpler (e.g., the income next year is close to the income from the previous year). Thus, we can build a realistic simulator by fitting the causal mechanisms of initial states to data and hand-designing transition mechanisms. The effects of interventions in early time steps on outcomes later in time are determined by compositions of several transitions which renders causal effects non-trivial functions of the initial state, even if each transition is simple. We use our simulator IncomeSCM to generate data for an observational study of the effects of studies on future income and compare popular estimators from the literature. Our results indicate that different estimators yield estimates that differ substantially in quality despite fitting observed variables similarly well, pointing to the challenging nature of the tasks.
@@ 2 Related work
Recognizing the limitations of relying on purely synthetic simulators for benchmarking, several works have proposed fully or partially data-driven simulators, where components are learned from real-world data. A recent example is Medkit-Learn Chan et al. (2021), developed for off-policy evaluation of reinforcement learning, and the Alzheimer’s Disease Causal estimation Benchmark (ADCB) (Kinyanjui and Johansson, 2022). The 2022 iteration of the ACIC challenge included a partially data-driven design, which used a strategy similar to ours, but to the best of our knowledge, the details of the simulator remain unpublished (Mathematica, 2022). Complementing purpose-built data sets and simulators, causality toolboxes, such as DoWhy (Sharma and Kiciman, 2020) and CausalML (Chen et al., 2020) include multiple ways of generating simple forms of synthetic data. Gentzel et al. (2019) provide an excellent overview of different approaches to evaluating causal modeling methods and suggest another approach: Use complex systems, such as a PostGres database, where interventions on its configuration can be performed but the outcome are hard to predict. This can provide both observational data for learning and interventional data for evaluation. Appendix D contains a tabular summary of common causal effect estimation benchmarks.
@@ 3 Constructing the simulator
IncomeSCM is based on a discrete-time Markov structural causal model (SCM). The primary goal of the model is to simulate the effects of an intervention of interest $A$ on a given outcome $Y$ in a system of random variables $\mathcal{V}\coloneqq\{A,Y\}\cup{\bf X}$, where ${\bf X}$ is a set of context variables. Each variable $V\in\mathcal{V}$ is simulated at $T$ time points in a sequence $V_{1},...,V_{T}$. The initial state of $V_{1}$ is determined by a deterministic structural equation $V_{1}:=f_{V}(\mathrm{\textbf{pa}}(V_{1}),U_{V_{1}})$ where $\mathrm{\textbf{pa}}(V_{1})$ are the parents (direct causes) of $V_{1}$ and $U_{V_{1}}$ is a source of noise. Values of $V_{t}$ for $t>1$ are determined by a time-homogenous transition mechanism $V_{t}:=g_{V}(\mathrm{\textbf{pa}}(V_{t}),U_{V_{t}})$. In general, mechanisms and parent sets are not shared by initial states and transitions, i.e., $f_{V}\neq g_{V}$ and $\mathrm{\textbf{pa}}(V_{1})\neq\mathrm{\textbf{pa}}(V_{t})=\mathrm{\textbf{pa}}(V_{s})$ for $s,t>1$. One of the main reasons for this is that transitions $g_{V}$ typically depend on the previous value of the variable $V$, but the initial state cannot. As a result, the initial state of a variable typically has a much stronger dependence on other variables in the systems than the transitions do. The model is Markov in that $\mathrm{\textbf{pa}}(V_{t})$ includes only variables from time $t$ and $t-1$.
@@
On hand-crafted transitions.
The design strategy for our simulator is developed for systems with simpler transitions than initial states. A strong argument for this idea is that most variables are easier to model once you know a recent value of them. The classical example is that the weather tomorrow is usually similar to the weather today (transition is simple), but modeling the weather based on other variables, such as location and time of year, is more difficult (the initial-state model is complex). Another example is the distribution of patients in a hospital. When a new patient gets admitted (initial state), we can view their demographics, symptoms, and medications as drawn from a complex distribution of possible patients. When we observe them the next day, the changes in their variable tend to be more predictable given their values at admission and obey simpler rules, such as dose response to medication or day-night cycles of vitals. A counter-example to this pattern would be a system where all variables are independent in the initial state but become dependent in transitions. For example, let the variables measure the opinions of a random set of people (e.g., students) who begin to meet regularly (say, in a class). In the initial state, there will be no correlation among the opinions but he next state may be a complicated function of the opinions of different people.
@@ 3.1 The base data set: Adult
As the primary outcome of interest $Y$, we use the yearly income of subjects in US dollars. In the Adult data set, income is coded as a binary variable, representing whether the yearly income of a subject exceeds \$50 000, denoted here by $R\in\{0,1\}$. For IncomeSCM, we construct a continuous income variable $Y$ by first fitting a small random forest regressor $h_{R}$ to predict the original binary outcome $R$ as a function of all covariates $X$ (excluding the added variable studies). The continuous predictions $\hat{R}_{i}=h_{R}(X_{i})\in[0,1]$ of the regressor are computed for each subject $i$. This intermediate value represents the likelihood that the income exceeds \$50 000, which is then re-scaled and shifted by two constants so that the average of the resulting numbers, across the full cohort, matches the average US population annual salary (in USD). Leaving the outcome a binary variable would make it more difficult to design, for example, income transitions. With a continuous value, we can reason about the average increase in salary due to a pay raise. A binary-income model would be much more coarse-grained.
@@
For a causal estimation task to be realistic, the primary intervention of interest must be possible to implement in practice, for example, in an experiment (Hernán, 2019; Hernán et al., 2022). The Adult data set contains a small number of variables that could be directly intervened upon and serve as the primary intervention. Base variables age, sex, native-country, and race can be removed from consideration, as can capital-gain, capital-loss as these are outside of the control of any feasible experiment. occupation is feasible to change on a yearly basis, but most jobs are unavailable to most people. relationship and marital-status are possible to change, through e.g., divorce, but most people would not be willing to change these in order to optimize their income, let alone participate in an experiment to do so. workclass, occupation and hours-per-week are candidates that could be intervened upon a yearly basis. education is coded in Adult by the highest attained degree (e.g., Bachelor’s degree) or the number of years of study. However, in the time scale of the simulator (one observation per year, $t=1,...,T$), an intervention of, say $A_{t}\leftarrow\mathrm{Doctorate}$ is not feasible. Instead, to showcase the blend of data-driven and knowledge-guided design, we create an intervention on studies, a constructed variable, with values “Full-time studies”, “Day course”, “Evening course” and “No studies”, representing the study activity of a person during a year. The initial state of studies is determined by a logistic regression with manually selected coefficients for age, relationship, and education, and intercepts for each of the study types. studies directly affect the education level, current income and the propensity to study in the next year, and indirectly the future income of a subject.
@@ 3.2 The causal graph
We construct a cDAG for the Adult data set (the initial-state cDAG for IncomeSCM-1.0) by assigning direct causes to each variable. age, sex, race, and native-country were taken to be base variables without causal parents in the data set. education was deemed to be causal of workclass, which in turn causes occupation and marital-status. marital-status and relationship are closely linked, and the causal direction was chosen to be from the former to the latter. Finally, hours-per-week and capital-net were determined by the previous variables, and income is the furthest downstream, directly caused also by studies, the hand-designed variable. The full graph is illustrated in Figure 1 (middle-left). A very similar graph was used by Nabi and Shpitser (2018) except, in their case, marital-status was a direct cause of education, not the other way around.
@@ 3.3 The samplers
where $f_{V}$ predicts the value of $V$ given its parents—$f_{V}$ is a model of the conditional mean, $\mathbb{E}[V\mid\mathrm{\textbf{pa}}(V)]$—and $\epsilon_{V}$ is a source of homoskedastic noise (e.g., Gaussian). All structural equations for initial-state variables, except for studies, have observations $(V,\mathrm{\textbf{pa}}(V))$ available (income is a special case as described above). For these, $f_{V}$ is implemented by fitting a regression (e.g., a Random forest regressor) to the observed data, and the scale of $\epsilon_{V}$ is chosen proportional to the mean squared regression error. capital-net is an exception to this form, since the majority of capital gains/losses is 0. For this, we blend a classifier that predicts whether the value will be non-zero and a conditional regressor for the case where the value is non-zero.
For categorical variables $V\in\{1,...,k\}$, we sample from a categorical distribution given by the softmax function $\sigma$ applied to map $f_{V}$ from parent variables to logits for the categories of $V$. By the Gumbel-Max trick (Gumbel, 1954), we can write this as a structural equation
@@
Hand-crafted samplers.
For equations with no available data, such as for studies and all variable transitions, the samplers are designed by hand. All transition samplers are described in Appendix Table 7. For the intervention variable studies, the initial-state distribution and transitions are described above. The demographic variables native-countryt, sext and racet are constant, and aget increments deterministically by 1 each time step. For the initial state of the outcome variable income1 and the transitions $g_{V}$ of most variables $V\in\mathcal{V}$, we use a blend of hand-designed rules and fitted regressions. This allows us to generate complex dependencies between variables without observing a time series. For example, income1 is based on a regression of the original binary variable in the Adult data set, with hand-crafted modifications depending on studies and workclass. The transition sampler of relationshipt has a fixed probability $p_{\mathrm{stay}}$, that the person will stay in their relationship, relationshipt = relationshipt-1. With probability $1-p_{\mathrm{stay}}$, relationshipt is resampled from the same fitted model as the initial state, see Figure 1 (right). In this way, only the event that the relationship is changed is hand-designed, but what it is changed to depends on data. A similar strategy is used for all context variables except demographics and education.
@@ 4 Implementation & data set
Here, parents specify the direct causes of the initial state of the variable. sampler specifies the type and parameters of the sampler. In this case, the initial state of education is distributed according to a logistic regression of age, race, sex and native-country. seq_parents_curr specifies the direct causes of the transition probability from the current time step, and seq_parents_prev causes from the previous time step. seq_sampler specifies which function determines this probability. LogisticSampler is a general sampler, that outputs the probability of the given variable as a logistic-linear function of the direct causes specified in parents. General samplers can be reused for other data sets and are defined in samplers.py.
@@ 4.1 Benchmark data set
For this task, we create a cross-sectional observational data set by first sampling from the simulator at time $t=1,...,7$ where studies are selected by a “default” observational stochastic policy as defined by the samplers in IncomeSCM. Then, time $t=1$ is used to record the previous income level income_prev and studies studies_prev, prior to intervention. All other covariates are read off at time $t_{0}=2$. The binary intervention of interest is then performed at $t_{0}=2$, with $A_{2}=1$ representing Full-time studies and $A_{2}=0$ No studies. The outcome variable $Y_{T}=$incomeT is recorded at $T=7$. We use this to sample $n=50000$ independent and identically distributed observation sequences, each representing an individual. The sequences are drawn with an incrementing random seed, starting at $s=0$. First-order covariate statistics of covariates, interventions, and outcomes for this seed can be seen in Table 1. There are no unobserved confounders or temporal confounding in this benchmark task—all direct causes of the intervention are observed in the cross-sectional data set.
@@
Identification strategies for the causal effect of any variable in the simulator on any downstream variable can be derived from the edge list in Table 6 using graphical criteria like the backdoor criterion (BDC) (Pearl, 2009). In the benchmark data set, the observed intervention is performed at time $t=2$ with $t=1$ being the initial state. By the BDC, a minimal adjustment set for estimating the causal effect of intervening on studies2 is the set of direct causes of this variable. In IncomeSCM-1.0, studiest at times $t>0$ are directly caused by studiest-1, incomet-1, aget, sext,educationt, and relationshipt. [^2] In the implementation, the helper variable time is also passed to enable interventions at specific time steps. Regression adjustment of potential outcomes or propensity-based adjustment using these variables leads to the identification of causal effects. Expanding this set with any pre-intervention variables (e.g., other variables at times $t=1,2$) also results in valid adjustment sets.
@@
    CATE, conditioned only on education2 (not itself a valid adjustment set), represented as 16 numeric bins, using the full adjustment set from Task 1.
Task 3 is an example of when the conditioning set of the CATE differs from any valid adjustment set. That is, to estimate the CATE conditioned only on education2, it is not sufficient to adjust only for this variable; it requires a slightly unusual application of common causal effect estimators.
@@ 5 Experiments
We include regression estimators in the form of T-learners, which fit one model per potential outcome, and S-learners, which treat the intervention as an input to a single model, mixed in with adjustment variables (Künzel et al., 2019). For both, we compare multiple base estimators in ridge regression (Ridge), random forests (RF), and XGBoost (XGB). Second, we use two classical inverse-propensity weighting estimators, the Horvitz-Thompson (IPW) and Hayek or “Weighted” (IPW-W). Third, we include three double-machine learning (DML) estimators created by fitting regressions to the target $\frac{y_{i}-M_{Y}(x_{i})}{a_{i}-M_{A}(x_{i})}$, where $M_{Y},M_{A}$ model $\mathbb{E}[Y\mid X]$ and $p(A=1\mid X)$ respectively, using sample weights $w_{i}=(a_{i}-M_{A}(x_{i}))^{2}$. This stems from the residualization of Robinson (1988), as used by Chernozhukov et al. (2018). The three variants use different base estimators of the nuisance functions. DML (Mix) uses an XGBoost regressor for $M_{Y}$, logistic regression for $M_{A}$, and linear regression for the cate estimator. Finally, we include a nearest-neighbor matching estimator with a Euclidean metric. The IPW, matching, and Ridge-regression S-learner estimators return estimates only of the ATE, not the CATE. For a more detailed description of all estimators, see Appendix F.
Hyperparameters for all estimators (see Appendix Table 8) were selected based on 5-fold cross-validation of 20 uniformly sampled hyperparameter settings. The estimators were then refitted to the entirety of the observational training data. Producing all results, including fitting the simulator, sampling, estimating, and evaluating causal effects, took less than 1 hour on a 2023 M2 MacBook Pro. The results were replicated on a 2 x 16 core Intel(R) Xeon(R) Gold 6226R CPU @ 2.90GHz with minimal differences except for XGBoost which is known to yield different fits for different machines.